No. cos de find using the approxinate Simpsons rule value with of 6 do Strips.
Answer #1

T1/2 I loss dr , assume f(n) = Coss. interval [o /] into 6 take n=6, ie we will divide the Subintervals then length of each s$3(0-5) dr - 3 (( - 100 +25) dr = 3(03) - 10 q +257 +c. 3_527 +258 +C; where is integrating constant.